# Scientific Large Language Models: A Survey on Biological & Chemical Domains

### v1

@@ 4.3. Evaluation
Property Prediction.
Molecular property prediction stands as a pivotal challenge in computational chemistry, where the goal is to accurately predict the properties of molecules, such as solubility, lipophilicity, affinity, absorption, distribution, metabolism, excretion, toxicity, and biological activity, based on their chemical structure. In this domain, most encoder-only models are typically pre-trained on extensive and unlabeled datasets and subsequently fine-tuned using specific data for property prediction. In the traditional domain of property prediction, BERT-based models such as ChemBERTa (Chithrananda et al., 2020), MAT (Maziarka et al., 2020), Mol-BERT  (Li and Jiang, 2021) and MG-BERT (Zhang et al., 2021) are widely recognized. However, in cases of small sample problems, SMILES Transformer (Honda et al., 2019) has demonstrated superior performance.
@@
Molecule Generation.
Molecule Generation is an emerging and critical area in computational chemistry and drug design, harnessing the power of advanced computational techniques to create novel molecular structures. This field is particularly vital in the development of new pharmaceuticals and materials, where the ability to generate molecules with desired properties can greatly accelerate the discovery process. Molecule generation encompasses two primary approaches: template-based design and de novo design. The former involves modifying known molecules or scaffolds. This approach starts with existing molecular frameworks, which are then altered to enhance desirable properties or reduce unwanted effects. The de novo design refers to the creation of novel molecular structures from scratch, without relying on pre-existing templates. This method is particularly valuable for discovering unique compounds with potential therapeutic or material applications. It challenges computational models to explore vast chemical spaces, requiring sophisticated algorithms capable of predicting viable and effective molecular structures. To address this task, there are lots of typical Mol-LLMs such as MolGPT (Bagal et al., 2021), SMILES GPT (Adilov, 2021) and Taiga (Mazuz et al., 2023).
Evaluation Metric To quantify the performance of Mol-LLMs in property prediction, interaction prediction, and reaction prediction tasks, which primarily involve classification and regression, one can employ the standard metrics in machine learning, such as accuracy, F1-score, and correlation coefficient. For molecular generation tasks, their performance is evaluated based on their ability to produce diverse, realistic, and relevant molecules. We present several popular metrics as follows.
@@
[…]
@@
[…]
@@ 5. Protein Large Language Models
(Sec. 5.1),text width=4em [Encoder-only, text width= 5.5em [Sequence-only: ProteinBert (Brandes et al., 2022)   ProtTrans (Elnaggar et al., 2022)  ESM-1b (Rives et al., 2019)
ESM-2 (Lin et al., 2022)  PMLM  (He et al., 2021)   ProtFlash (Wang et al., 2023e)   ProteinNPT  (Notin et al., 2023b) ] [MSA-combined: ESM-MSA-1b (Rao et al., 2021)  ESM-1v (Meier et al., 2021)   RSA (Ma et al., 2023) ] [Add 3D structure: ESM-GearNet (Zhang et al., 2023d)  SaProt (Su et al., 2023)  LM-GVP (Wang et al., 2022a)  Mansoor et al. (Mansoor et al., 2021) ] [Integrating multi-level structure: PromptProtein (Wang et al., 2023d)  ] [Add extra knowledge: OntoProtein (Zhang et al., 2022)  KeAP (Zhou et al., 2023) ] ] [Decoder-only, text width= 5.5em [GPT-based: ProGen  (Madani et al., 2020)  ProGen2 (Nijkamp et al., 2023)   RITA (Hesslow et al., 2022)  ProtGPT2 (Ferruz et al., 2022)
LM-D Design (Zheng et al., 2023a)  PoET (au2 and Bepler, 2023)  IgLM (Shuai et al., 2021)  ZymCTRL (Munsamy et al., 2022) ] ] [Encoder-decoder, text width=5.5em [T5-based: ProstT5  (Heinzinger et al., 2023)  pAbT5  (Chu and Wei, 2023) ] [Transformer-based: SS-pLM (Serrano et al., 2023) ] [GLM-based: xTrimoPGLM (Chen et al., 2023a) ] [MSA-combined: MSA2Prot  (Ram and Bepler, 2022)  MSA-Augmenter  (Zhang et al., 2023)  EvoOpt (Yamaguchi and Saito, 2022)  Sgarbossa et al. (Sgarbossa et al., 2023) ] [Others: Fold2Seq (Cao et al., 2021)  Lee et al. (Lee et al., 2022) ] ] ] [Datasets
@@
ProteinGym (Notin et al., 2023a)  FLIP (Dallago et al., 2021)  PEER (Xu et al., 2022)  TAPE (Rao et al., [n. d.]) ] ] ] [Evaluation
@@ 5.1.1. Encoder-only Models
Most of the encoder-only Prot-LLMs are built upon the encoder of Transformer, which enables the encoding of protein sequences or structures into fixed-length vector representations. These representations facilitate the identification of patterns and features within proteins, thereby enhancing subsequent analysis and prediction tasks. Self-supervised learning is employed to acquire protein representation, such as masked language modeling (MLM) tasks that reconstruct corrupted tokens based on the surrounding sequence. Prominent pre-trained protein sequence encoders include ESM-1b (Rives et al., 2019), ESM-1v (Meier et al., 2021), and ESM-2 (Lin et al., 2022), ProteinBert (Brandes et al., 2022), ProtTrans (Elnaggar et al., 2022). The ESM series (Rives et al., 2019; Meier et al., 2021; Lin et al., 2022) mainly employs a Transformer’s encoder (i.e., BERT (Devlin et al., 2018) and RoBERTa (Liu et al., 2019)) to predict protein structure and function, which utilizes extensive sequence information from protein databases without relying on manual annotations of the sequences. ProteinBert (Brandes et al., 2022) enhances the classical BERT architecture by incorporating a novel pretraining task for predicting protein functionality. It separates local (character-level) and global (sequence-level) representations, enabling multitask processing in a principled way. ProtTrans (Elnaggar et al., 2022) trained several auto-encoder models (includes BERT (Devlin et al., 2018), Albert (Lan et al., 2019), Electra (Clark et al., 2020)) on a vast of sequence data. Several works have attempted to improve the encoder architecture and the training method. PMLM (He et al., 2021) introduced Pairwise Masked Language Model (PMLM), to directly pretrain the encoder with a focus on capturing co-evolutionary information reflected in residue co-variation within the sequence. This approach considers dependencies between masked tokens. ProtFlash (Wang et al., 2023e) suggested a mixed-chunk attention mechanism that combines multiple positional encodings and a hybrid block attention mechanism incorporating both local and global attention, effectively reducing model complexity. ProteinNPT (Notin et al., 2023b) developed a non-parametric Transformer specifically designed for protein sequences, making it particularly suitable for scenarios involving sparse labels and multi-task learning.
@@
Given that the function of a protein is intricately linked to its structure, another reasonable approach for protein representation would involve the integration of its 3D conformation. There are some recent efforts in this direction. ESM-GearNet (Zhang et al., 2023d) combines ESM-1b and GearNet (Xie et al., 2022) (the state-of-the-art protein structure encoder) to enhance model performance. SaProt (Su et al., 2023) introduces a structure-aware vocabulary that integrates residue tokens with structure tokens. LM-GVP (Wang et al., 2022a) compose Transformer blocks and a graph network derived from the protein 3D structure. Mansoor et al. (Mansoor et al., 2021) encoded protein sequence and structure through joint training in a semi-supervised manner. PromptProtein (Wang et al., 2023d) utilizes prompt-guided multi-task pretraining to fuse different levels of protein structure. It is worth noting that numerous 3D structure encoders (like GearNet (Xie et al., 2022) and GBPNet (Aykent and Xia, 2022)) have been proposed to model protein structural information; however, they are beyond the scope of this survey which primarily focuses on LLMs.
Additionally, some approaches propose to introduce extra knowledge such as Gene Ontology (GO) to enhance the protein representation. OntoProtein (Zhang et al., 2022) considered GO as a factual knowledge graph, which was used to enhance protein representations. KeAP (Zhou et al., 2023) proposed a knowledge-injected auto-encoder that performs token-level knowledge graph exploration for protein representation learning.
@@
PMLM (He et al., 2021) 2021.07 87M - 731M Trans. enc. Uniref50/Pfam Contact pred. $\times$
@@
PromptProtein  (Wang et al., 2023d) 2023.02 650M RoBERTa UniRef50, PDB Func. pred.
@@
ProtFlash  (Wang et al., 2023e) 2023.10 79M/174M Trans. enc UniRef50 Func. pred.
ESM-GearNet  (Zhang et al., 2023d) 2023.10 - ESM-1b, GearNet - Func. pred.
@@
xTrimoPGLM  (Chen et al., 2023a) 2023.07 100B GLM Uniref90, ColdFoldDB Prot. gen., Func. pred. $\times$
@@ 5.1.2. Decoder-only Models
The decoder-based protein language models have been widely applied in specific protein designs. Timothy et al. introduced the protein evolutionary Transformer (PoET), an autoregressive generative model of the distribution over protein families, which can generate sets of related proteins. Zheng et al. proposed the LM-Design (Zheng et al., 2023a) that uses language models for structure-based protein design (i.e., inverse folding). ZymCTRL (Munsamy et al., 2022) trained a conditional language model on the BRENDA enzyme database (Chang et al., 2021), aimed at designing customized artificial enzymes by generating specific enzyme classes based on user prompts. IgLM (Shuai et al., 2021) employed autoregressive sequence generation for antibody design.
@@ 5.1.3. Encoder-decoder Models
The encoder-decoder architecture is commonly used for sequence-to-sequence tasks. In the field of protein research, ProstT5  (Heinzinger et al., 2023), a protein language model based on the T5 architecture (Raffel et al., 2020), is an exemplary encoder-decoder model that facilitates translation between protein sequences and structures. The conversion of protein structure from 3D to 1D is achieved through the utilization of 3Di-tokens introduced by Foldseek (van Kempen et al., 2022). pAbT5 (Chu and Wei, 2023) is another T5-based model that takes into consideration the constraints imposed by protein-protein interactions on generation and design. It can generate complementary heavy or light chains from their pairing partners. xTrimoPGLM (Chen et al., 2023a), based on General Language Model (GLM) architecture (Du et al., 2021), is a unified protein language model for understanding and generating protein sequences. It has an unprecedented scale of 100B parameters and consumes 1 trillion training tokens. Considering that the training of protein LLMs involves millions to billions of sequences and billions of parameters, Serrano et al. (Serrano et al., 2023) introduced the Small-Scale protein Language Model (SS-pLM), which significantly reduces the computational burden, democratizing the use of foundational models in protein generation.
@@
In the field of structure-based protein design, Cao et al. (Cao et al., 2021) proposed a Fold2Seq framework, a generative approach based on the Transformer architecture, which aims at designing protein sequences conditioned on specific target folds. To capture intricate sequence-structure relationships, Fold2Seq jointly learns embeddings for both sequences and folding structures, which are derived from the density of secondary structure elements in 3D voxels. Additionally, some approaches utilize reinforcement learning to improve the generation quality. Lee et al. (Lee et al., 2022) efficiently navigates the latent representation space rather than the protein sequence space by formulating the sequence design task as a Markov decision process and employing model-based reinforcement learning. This approach enables them to generate proteins with enhanced functionality and cellular fitness.
@@
Dataset Time Scale Keywords
@@
TAPE (Rao et al., [n. d.]) 2021.09 $\sim 120$K Remote homology detection, Secondary structure, Contact, Fluorescence, Stability prediction

### v2

@@ 4.3. Evaluation
Property Prediction. Molecular property prediction stands as a pivotal challenge in computational chemistry, where the goal is to accurately predict the properties of molecules, such as solubility, lipophilicity, affinity, absorption, distribution, metabolism, excretion, toxicity, and biological activity, based on their chemical structure. In this domain, most encoder-only models are typically pre-trained on extensive and unlabeled datasets and subsequently fine-tuned using specific data for property prediction. In the traditional domain of property prediction, BERT-based models such as ChemBERTa (Chithrananda et al., 2020), MAT (Maziarka et al., 2020), Mol-BERT  (Li and Jiang, 2021) and MG-BERT (Zhang et al., 2021) are widely recognized. However, in cases of small sample problems, SMILES Transformer (Honda et al., 2019) has demonstrated superior performance.
@@
Molecule Generation. Molecule Generation is an emerging and critical area in computational chemistry and drug design, harnessing the power of advanced computational techniques to create novel molecular structures. This field is particularly vital in the development of new pharmaceuticals and materials, where the ability to generate molecules with desired properties can greatly accelerate the discovery process. Molecule generation encompasses two primary approaches: template-based design and de novo design. The former involves modifying known molecules or scaffolds. This approach starts with existing molecular frameworks, which are then altered to enhance desirable properties or reduce unwanted effects. The de novo design refers to the creation of novel molecular structures from scratch, without relying on pre-existing templates. This method is particularly valuable for discovering unique compounds with potential therapeutic or material applications. It challenges computational models to explore vast chemical spaces, requiring sophisticated algorithms capable of predicting viable and effective molecular structures. To address this task, there are lots of typical Mol-LLMs such as MolGPT (Bagal et al., 2021), SMILES GPT (Adilov, 2021) and Taiga (Mazuz et al., 2023).
Evaluation Metric. To quantify the performance of Mol-LLMs in property prediction, interaction prediction, and reaction prediction tasks, which primarily involve classification and regression, one can employ the standard metrics in machine learning, such as accuracy, F1-score, and correlation coefficient. For molecular generation tasks, their performance is evaluated based on their ability to produce diverse, realistic, and relevant molecules. We present several popular metrics as follows.
@@
  • Internal diversity (IntDivp) (Polykovskiy et al., 2020). This metric ensures that the molecules produced by generative models exhibit a wide range of chemical properties, which is defined by
@@
    The formula IntDivp($G$) evaluates the diversity of a molecule set $G$, where $G$ is the set of molecules, $m_{1}$ and $m_{2}$ are individual molecules in $G$, $T$ is the similarity measure between two molecules, and $p$ adjusts the sensitivity to similarity.
@@ 5. Protein Large Language Models
(Sec. 5.1),text width=4em [Encoder-only, text width= 5.5em [Sequence-only: ProteinBert (Brandes et al., 2022)  ProtTrans (Elnaggar et al., 2022)  ESM-1b (Rives et al., 2019)  ESM-2 (Lin et al., 2022)
PMLM  (He et al., 2021b)  ProtFlash (Wang et al., 2023i)   ProteinNPT  (Notin et al., 2023b)  ] [MSA-combined: ESM-MSA-1b (Rao et al., 2021)  ESM-1v (Meier et al., 2021)  RSA (Ma et al., 2023) ] [Add 3D structure: ESM-GearNet (Zhang et al., 2023e)  SaProt (Su et al., 2023)  LM-GVP (Wang et al., 2022a)
Mansoor et al. (Mansoor et al., 2021)  ESM3 (Hayes et al., 2024) ] [Integrating multi-level structure: PromptProtein (Wang et al., 2023h)  ] [Add extra knowledge: OntoProtein (Zhang et al., 2022)  ESM All-Atom  (Zheng et al., 2024)  KeAP (Zhou et al., 2023)
KnowRLM  (Wang et al., [n. d.])  Outeiral et al.  (Outeiral and Deane, 2024)  ] ] [Decoder-only, text width= 5.5em [GPT-based: ProGen  (Madani et al., 2020)  ProGen2 (Nijkamp et al., 2023)   RITA (Hesslow et al., 2022)  ProtGPT2 (Ferruz et al., 2022)
PoET (au2 and Bepler, 2023)  IgLM (Shuai et al., 2021)  ZymCTRL (Munsamy et al., 2022)  ProteinRL (Sternke and Karpiak, 2023)  C. Frey et al.  (Frey et al., 2023)  ] ] [Encoder-decoder, text width=5.5em [T5-based: ProstT5  (Heinzinger et al., 2023)  pAbT5  (Chu and Wei, 2023) ] [Transformer-based: SS-pLM (Serrano et al., 2023)  ESM-GearNet-INR-MC  (Lee et al., 2023) ] [GLM-based: xTrimoPGLM (Chen et al., 2023b) ] [MSA-combined: MSA2Prot  (Ram and Bepler, 2022)  MSA-Augmenter  (Zhang et al., 2023)  EvoOpt (Yamaguchi and Saito, 2022)  Sgarbossa et al. (Sgarbossa et al., 2023) ] [Others: Fold2Seq (Cao et al., 2021)  Lee et al. (Lee et al., 2022)  LM-Design (Zheng et al., 2023a) ] ] ] [Datasets
@@
ProteinGym (Notin et al., 2023a)  FLIP (Dallago et al., 2021)  PEER (Xu et al., 2022)  TAPE (Rao et al., 2019) ] ] ] [Evaluation
@@ 5.1.1. Encoder-only Models
Most of the encoder-only Prot-LLMs are built upon the encoder of Transformer, which enables the encoding of protein sequences or structures into fixed-length vector representations. These representations facilitate the identification of patterns and features within proteins, thereby enhancing subsequent analysis and prediction tasks. Self-supervised learning is employed to acquire protein representation, such as masked language modeling (MLM) tasks that reconstruct corrupted tokens based on the surrounding sequence. Prominent pre-trained protein sequence encoders include ESM-1b (Rives et al., 2019), ESM-1v (Meier et al., 2021), and ESM-2 (Lin et al., 2022), ProteinBert (Brandes et al., 2022), ProtTrans (Elnaggar et al., 2022). The ESM series (Rives et al., 2019; Meier et al., 2021; Lin et al., 2022) mainly employs a Transformer’s encoder (i.e., BERT (Devlin et al., 2018) and RoBERTa (Liu et al., 2019)) to predict protein structure and function, which utilizes extensive sequence information from protein databases without relying on manual annotations of the sequences. ProteinBert (Brandes et al., 2022) enhances the classical BERT architecture by incorporating a novel pretraining task for predicting protein functionality. It separates local (character-level) and global (sequence-level) representations, enabling multitask processing in a principled way. ProtTrans (Elnaggar et al., 2022) trained several auto-encoder models (includes BERT (Devlin et al., 2018), Albert (Lan et al., 2019), Electra (Clark et al., 2020)) on a vast of sequence data. Several works have attempted to improve the encoder architecture and the training method. PMLM (He et al., 2021b) introduced Pairwise Masked Language Model (PMLM), to directly pretrain the encoder with a focus on capturing co-evolutionary information reflected in residue co-variation within the sequence. This approach considers dependencies between masked tokens. ProtFlash (Wang et al., 2023i) suggested a mixed-chunk attention mechanism that combines multiple positional encodings and a hybrid block attention mechanism incorporating both local and global attention, effectively reducing model complexity. ProteinNPT (Notin et al., 2023b) developed a non-parametric Transformer specifically designed for protein sequences, making it particularly suitable for scenarios involving sparse labels and multi-task learning.
@@
Given that the function of a protein is intricately linked to its structure, another reasonable approach for protein representation would involve the integration of its 3D conformation. There are some recent efforts in this direction. ESM-GearNet (Zhang et al., 2023e) combines ESM-1b and GearNet (Xie et al., 2022) (the state-of-the-art protein structure encoder) to enhance model performance. SaProt (Su et al., 2023) introduces a structure-aware vocabulary that integrates residue tokens with structure tokens. LM-GVP (Wang et al., 2022a) compose Transformer blocks and a graph network derived from the protein 3D structure. Mansoor et al. (Mansoor et al., 2021) encoded protein sequence and structure through joint training in a semi-supervised manner. PromptProtein (Wang et al., 2023h) utilizes prompt-guided multi-task pretraining to fuse different levels of protein structure. ESM3 (Hayes et al., 2024) is an advanced multimodal generative language model that effectively integrates protein sequence, structure, and function analysis. It exhibits exceptional capability in comprehending intricate prompts by leveraging its evolutionary simulation capabilities, making it highly adaptable for biological alignment tasks. It is worth noting that numerous 3D structure encoders (like GearNet (Xie et al., 2022) and GBPNet (Aykent and Xia, 2022)) have been proposed to model protein structural information; however, they are beyond the scope of this survey which primarily focuses on LLMs.
Additionally, some approaches propose to introduce extra knowledge such as Gene Ontology (GO) to enhance the protein representation. OntoProtein (Zhang et al., 2022) considered GO as a factual knowledge graph, which was used to enhance protein representations. ESM All-Atom (Zheng et al., 2024) achieves atom-scale and residue-scale unified molecular modeling by pretraining on multi-scale protein sequences and utilizing a multi-scale position encoding to capture relationships among residues and atoms. KeAP (Zhou et al., 2023) proposed a knowledge-injected auto-encoder that performs token-level knowledge graph exploration for protein representation learning. KnowRLM  (Wang et al., [n. d.]) stands out for its ability to utilize contextual amino acid information from knowledge graphs, thus attaining advantages from both statistical patterns of protein sequences and biochemical properties of amino acids. Some studies improve protein learning representation by enriching pre-trained biological data. For example, Outeiral et al. (Outeiral and Deane, 2024) show that large language models trained on codons, instead of amino acid sequences, provide high-quality representations that outperform comparable state-of-the-art models across a variety of tasks.
@@
PMLM (He et al., 2021b) 2021.07 87M - 731M Trans. enc. Uniref50/Pfam Contact pred. $\times$
@@
PromptProtein  (Wang et al., 2023h) 2023.02 650M RoBERTa UniRef50, PDB Func. pred.
@@
ProtFlash  (Wang et al., 2023i) 2023.10 79M/174M Trans. enc UniRef50 Func. pred.
ESM-GearNet  (Zhang et al., 2023e) 2023.10 - ESM-1b, GearNet - Func. pred.
@@
Outeiral et al.  (Outeiral and Deane, 2024) 2024.02 10M - 5B Trans. enc. European Nucleotide Archive Protein represent learning
ESM All-Atom  (Zheng et al., 2024) 2024.06 35M RoBERTa AlphaFold DB Unified Molecular Modeling $\times$
KnowRLM  (Wang et al., [n. d.]) 2024.06 - Trans. enc. - Protein Directed Evolution $\times$
ESM3  (Hayes et al., 2024) 2024.06 98B RoBERTa PDB Seq. pred., Func. pred., Struct. pred.
@@
ProteinRL (Sternke and Karpiak, 2023) 2023.10 764M GPT - Prot. design $\times$
@@
C. Frey et al.  (Frey et al., 2023) 2024.03 9.87M/1.03M GPT hu4D5 antibody mutant Functional prot. gen. $\times$
@@
xTrimoPGLM  (Chen et al., 2023b) 2023.07 100B GLM Uniref90, ColdFoldDB Prot. gen., Func. pred. $\times$
@@
ESM-GearNet- INR-MC (Lee et al., 2023) 2024.04 - Transformer Swiss-Prot, AlphaFoldDB Prot. gen $\times$
@@ 5.1.2. Decoder-only Models
The decoder-based protein language models have been widely applied in specific protein designs. Timothy et al. introduced the protein evolutionary Transformer (PoET), an autoregressive generative model of the distribution over protein families, which can generate sets of related proteins. Zheng et al. proposed the LM-Design (Zheng et al., 2023a) that uses language models for structure-based protein design (i.e., inverse folding). ZymCTRL (Munsamy et al., 2022) trained a conditional language model on the BRENDA enzyme database (Chang et al., 2021), aimed at designing customized artificial enzymes by generating specific enzyme classes based on user prompts. IgLM (Shuai et al., 2021) employed autoregressive sequence generation for antibody design. C. Frey et al. (Frey et al., 2023) evaluate the robustness of our approach on generative modeling of antibody proteins and introduce the distributional conformity score to benchmark protein generative models. ProteinRL (Sternke and Karpiak, 2023) is a reinforcement learning framework for fine-tuning generative PLMs for proteins optimized for specific sequence and/or structural properties. ProteinRL fine-tuning guides the PLM towards generating sequences optimized for the defined properties, extending to values rarely or never seen in natural sequences or sequences generated without ProteinRL fine-tuning.
@@ 5.1.3. Encoder-decoder Models
The encoder-decoder architecture is commonly used for sequence-to-sequence tasks. In the field of protein research, ProstT5  (Heinzinger et al., 2023), a protein language model based on the T5 architecture (Raffel et al., 2020), is an exemplary encoder-decoder model that facilitates translation between protein sequences and structures. The conversion of protein structure from 3D to 1D is achieved through the utilization of 3Di-tokens introduced by Foldseek (van Kempen et al., 2022). pAbT5 (Chu and Wei, 2023) is another T5-based model that takes into consideration the constraints imposed by protein-protein interactions on generation and design. It can generate complementary heavy or light chains from their pairing partners. xTrimoPGLM (Chen et al., 2023b), based on General Language Model (GLM) architecture (Du et al., 2021), is a unified protein language model for understanding and generating protein sequences. It has an unprecedented scale of 100B parameters and consumes 1 trillion training tokens. Considering that the training of protein LLMs involves millions to billions of sequences and billions of parameters, Serrano et al. (Serrano et al., 2023) introduced the Small-Scale protein Language Model (SS-pLM), which significantly reduces the computational burden, democratizing the use of foundational models in protein generation.
@@
In the field of structure-based protein design, Cao et al. (Cao et al., 2021) proposed a Fold2Seq framework, a generative approach based on the Transformer architecture, which aims at designing protein sequences conditioned on specific target folds. To capture intricate sequence-structure relationships, Fold2Seq jointly learns embeddings for both sequences and folding structures, which are derived from the density of secondary structure elements in 3D voxels.
While recent studies have successfully employed sequence- or structure-based representations to address multiple tasks in protein science, there has been significant oversight in incorporating protein surface information, a critical factor for protein function. Therefore, ESM-GearNet-INR-MC  (Lee et al., 2023) presents a pre-training strategy that incorporates information from protein sequences, 3D structures, and surfaces to improve protein representation learning.
Additionally, some approaches utilize reinforcement learning to improve the generation quality. Lee et al. (Lee et al., 2022) efficiently navigates the latent representation space rather than the protein sequence space by formulating the sequence design task as a Markov decision process and employing model-based reinforcement learning. This approach enables them to generate proteins with enhanced functionality and cellular fitness.
@@
Dataset Last updated Scale Keywords
@@
TAPE (Rao et al., 2019) 2021.09 $\sim 120$K Remote homology detection, Secondary structure, Contact, Fluorescence, Stability prediction
Reactome (Jassal et al., 2020) 2023.12 $\sim 3$M Biological interactions and pathways
STRING (Szklarczyk et al., 2023) 2022.11 59.3M Protein-Protein interaction networks
BioGRID (Oughtred et al., 2019) 2023.12 271k Genetic and protein interactions
InterPro (Paysan-Lafosse et al., 2023) 2024.01 $\sim 41$k Classification of protein families
DIP (Xenarios et al., 2002) 2020.02 81K Protein-protein interaction information
PROSITE (Sigrist et al., 2012) 2024.03 103K Collection of signatures that identify patterns or profiles in proteins